17. Suppose A is the 3x 3 zero matrix (i.e., all the entries are 0). Describe the solution set of the equation
Answer #1

Solution:

\small AX=0

0001 Га 0001 1b1=10 000 c

Since the rank of \small A   is \small 0< 3 , the number of unknown

\small \therefore The number of free variables are   \small 3-0=3

\small \therefore 0\cdot x+0\cdot y+0\cdot z=0

Put \small x=s\: ,\: y=t\: ,\: c=r

\small \therefore X=\left \{ \begin{bmatrix} s\\ t\\ r\\ \end{bmatrix}=s\begin{bmatrix} 1\\ 0\\ 0\\ \end{bmatrix}+t\begin{bmatrix} 0\\ 1\\ 0\\ \end{bmatrix}+r\begin{bmatrix} 0\\ 0\\ 1\\ \end{bmatrix}:s\: ,t\:, r\in \mathbb{R} \right \}

\small i.e\: \: A has infinitely many solution.